Session: Oct 2011 (session 3)
Wavelength: 1.3 cm
To give feedback on this experiment, click
Hartebeesthoek (Jonathan Quick at Thu Nov 10 05:42:29 2011) Success:
JIVE (Miguel Sanz at Thu Nov 10 05:44:47 2011) Success:
No know problems
(Michael Lindqvist at Thu Nov 10 07:33:42 2011) Success:
No known problems.
(Michael Lindqvist at Thu Nov 10 07:33:55 2011) Success:
No known problems. Light rain.
Yebes (Miguel Sanz at Thu Nov 10 08:18:38 2011) Success:
Sorry for using JIVE site instead Yebes.
(Wenjun Yang at Thu Nov 10 16:15:23 2011) Minor failures: HIGH
Effelsberg (Uwe Bach at Mon Nov 14 09:20:43 2011) Success:
Effelsberg (Uwe Bach at Mon Nov 14 09:21:51 2011) Success:
Effelsberg (Uwe Bach at Mon Nov 14 09:24:01 2011) Success:
and now with comment: Tsys for the BBC on Ori KL increased to 4000 K, but did not overflow.
Metsahovi (Rastorgueva at Mon Nov 14 14:56:19 2011) Success:
The session was carried out without any noticebla problems